# thanks: – corresponding author

### v2

@@ 3.2 Physical model realization
where $\omega_{1}$ and $\omega_{2}$ are the characteristic frequencies of the simple harmonic oscillator along the $x_{1}$- and $x_{2}$-direction, respectively. Simultaneously, let us recall (cf. the discussion in Sec.1) that the squeezing of bipartite coherent states can be obtained by acting the bipartite squeezing operators $\hat{S}_{k}(1,2;\xi)$ and the translation operators $\hat{D}(1,2;z_{1},z_{2})$ on the tensor product of two single-mode vacuum states. In consequence, we can expect that the eigenequation corresponding to the Hamiltonian (53), i.e.
@@
The Hamiltonians $\hat{H}_{k}(1,2;z_{1},z_{2},\xi)$ obtained in this manner still describe systems of the anisotropic harmonic oscillators, albeit with the requisite modifications. Therefore, it seems reasonable to presume that the construction of these Hamiltonians is based on the assumption that there exists a pair of the oscillator annihilation operators $\hat{c}_{1}$ and $\hat{c}_{2}$, and their conjugate creation operators $\hat{c}^{\dagger}_{1}$ and $\hat{c}^{\dagger}_{2}$, which are elements of an algebra based on the canonical commutation relations: $[\hat{c}_{i},\hat{c}_{j}^{\dagger}]=\hat{1}\delta_{ij}$, with $[\hat{c}_{i},\hat{c}_{j}]=\hat{0}$, where $i,j=1,2$. According to this statement, we can assume that the annihilation and the creation operators are given by
@@
where they satisfy the new cannonical commutation relations in the form: $[\hat{C}_{k;i},\hat{C}_{k;j}^{\dagger}]=(\hat{1}\delta_{ij})_{k}$, $[\hat{C}_{k;i},\hat{C}_{k;j}]=\hat{0}$, $[\hat{C}_{k;i}^{\dagger},\hat{C}_{k;j}^{\dagger}]=\hat{0}$, for $i,j=1,2$ and fixed $k$. In general, each of the operators expressed by the formulas (61)–(64) can be transformed into the form
@@ 4 Conclusions
Following the current consensus in quantum optics, the bipartite squeezed coherent states are generated by applying the squeezing operator acting on the two-mode vacuum state. In the Bargmann space, this construction may be realized in at least two ways, depending on the choice of linear combination coefficients with respect to the established basis. In this study, we consider two sets of linear combination coefficients: one comprising the products of the holomorphic Hermite polynomials of one complex variable, and the other is formed from the holomorphic Hermite polynomials of two complex variables. These coefficients can be regarded as the result of acting on the unimodal or bipartite realizations of the $\mathfrak{su}(1,1)$ algebra on the basis formed from the products of monomials in the Bargmann space. By applying the Segal-Bargmann transform, we go to the Schrödinger space, wherein the wave functions corresponding to the bipartite squeezed coherent states are expressed by the Gaussian functions. These functions take a factorizable or non-factorizable form depending on the method of squeezing employed. This naturally gives rise to the question of whether the states thus obtained are entangled. To address this question, we have employed the Horodecki-Peres criterion for continuous-variable systems adapted to the phase-space approach. As demonstrated by Simon, the realization of this criterion in the phase space has a straightforward interpretation in terms of the Wigner distribution functions. This is because it has been sufficient to limit our examination to determine the symplectic eigenvalues of the appropriately modified covariance matrix of the Gaussians related to the Wigner distribution functions, following Simon’s recipe. The result we have obtained explicitly depends on the value of the squeezing parameter, $\xi=-(\log{\alpha})/2$, where the parameter $\alpha\in(0,1)$. Furthermore, we have demonstrated that applying the quantitative log-negativity criterion for the entanglement allows us to state that the no squeezing $\xi=0$ is equivalent to the absence of the entanglement in the examined states. Conversely, when the squeezing parameter $\xi\neq 0$, the states are entangled, wherein the maximal entanglement, exemplified by the EPR states, occurs when the parameter value approaches infinity.
In order to demonstrate the findings of our study, we have proposed and investigated a theoretical model of a two-particle mechanical system that allows expounding physical interpretation of the presented result. The Hamiltonian we have derived describe a system consisting of two coupled anisotropic harmonic oscillators depending on the parameter $\alpha$. We have demonstrated that the limit form that corresponds to $\alpha\rightarrow 1_{-}$ or alternatively $\xi\to 0_{+}$ of the aforementioned Hamiltonian reduces to the Hamiltonian of the two independent anisotropic harmonic oscillators shifted in positions and momenta. Furthermore, we also have shown that assuming no shifts and assuming identical oscillation frequencies, the Hamiltonian we have obtained is consistent with the simplified model discussed in the previous work.
@@ Appendix A Appendix: Mehler formulae
$\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}\frac{s^{m}t^{n}}{m!n!}H_{m,n}(z_{1},_{2})=\exp(sz_{1}+tz_{2}-st)$ (88)
@@
$=$ $e^{-\frac{1}{4}(\partial_{u}^{2}+\partial_{v}^{2})}\exp{\left[\sqrt{2}usz_{1}+\sqrt{2}vtz_{2}-2usvt\right]}$

### v3

@@ III.2 Physical model realization
where $\omega_{1}$ and $\omega_{2}$ are the characteristic frequencies of the simple harmonic oscillator along the $x_{1}$- and $x_{2}$-direction, respectively. Simultaneously, let us recall (cf. the discussion in Sec.I) that the squeezing of bipartite coherent states can be obtained by acting the bipartite squeezing operators $\hat{S}_{k}(1,2;\xi)$ and the translation operators $\hat{D}(1,2;z_{1},z_{2})$ on the tensor product of two single-mode vacuum states. In consequence, we can expect that the eigenequation corresponding to the Hamiltonian (53), i.e.
@@
The Hamiltonians $\hat{H}_{k}(1,2;z_{1},z_{2},\xi)$ obtained in this manner still describe systems of the anisotropic harmonic oscillators, albeit with the requisite modifications. Therefore, it seems reasonable to presume that the construction of these Hamiltonians is based on the assumption that there exists a pair of the oscillator annihilation operators $\hat{c}_{1}$ and $\hat{c}_{2}$, and their conjugate creation operators $\hat{c}^{\dagger}_{1}$ and $\hat{c}^{\dagger}_{2}$, which are elements of an algebra based on the canonical commutation relations: $\hat{c}_{i},\hat{c}_{j}^{\dagger}]=\hat{1}\delta_{ij}$, with $\hat{c}_{i},\hat{c}_{j}]=\hat{0}$, where $i,j=1,2$. According to this statement, we can assume that the annihilation and the creation operators are given by
@@
where they satisfy the new cannonical commutation relations in the form: $\hat{C}_{k;i},\hat{C}_{k;j}^{\dagger}]=(\hat{1}\delta_{ij})_{k}$, $\hat{C}_{k;i},\hat{C}_{k;j}]=\hat{0}$, $\hat{C}_{k;i}^{\dagger},\hat{C}_{k;j}^{\dagger}]=\hat{0}$, for $i,j=1,2$ and fixed $k$. In general, each of the operators expressed by the formulas (61)–(64) can be transformed into the form
@@ IV Conclusions
Following the current consensus in quantum optics, the bipartite squeezed coherent states are generated by applying the squeezing operator acting on the two-mode vacuum state. In the Bargmann space, this construction may be realized in at least two ways, depending on the choice of linear combination coefficients with respect to the established basis. In this study, we consider two sets of linear combination coefficients: one comprising the products of the holomorphic Hermite polynomials of one complex variable, and the other is formed from the holomorphic Hermite polynomials of two complex variables. These coefficients can be regarded as the result of acting on the unimodal or bipartite realizations of the $\mathfrak{su}(1,1)$ algebra on the basis formed from the products of monomials in the Bargmann space. By applying the Segal-Bargmann transform, we go to the Schrödinger space, wherein the wave functions corresponding to the bipartite squeezed coherent states are expressed by the Gaussian functions. These functions take a factorizable or non-factorizable form depending on the method of squeezing employed. This naturally gives rise to the question of whether the states thus obtained are entangled. To address this question, we have employed the Horodecki-Peres criterion for continuous-variable systems adapted to the phase-space approach. As demonstrated by Simon, the realization of this criterion in the phase space has a straightforward interpretation in terms of the Wigner distribution functions. This is because it has been sufficient to limit our examination to determine the symplectic eigenvalues of the appropriately modified covariance matrix of the Gaussians related to the Wigner distribution functions, following Simon’s recipe. The result we have obtained explicitly depends on the value of the squeezing parameter, $\xi=-(\log{\alpha})/2$, where the parameter $\alpha\in(0,1)$. Furthermore, we have demonstrated that applying the quantitative log-negativity criterion for the entanglement allows us to state that the no squeezing $\xi=0$ is equivalent to the absence of the entanglement in the examined states. Conversely, when the squeezing parameter $\xi\neq 0$, the states are entangled, wherein the maximal entanglement, exemplified by the EPR states, occurs when the parameter value approaches infinity. One of the central results of our findings presented in this paper is the following observation. For the families of bipartite Gaussian functions considered in this paper, we conclude, based on the Simon-Horodecki-Peres criterion, that the considered Gaussian wave functions expressed in the originally defined position coordinates, denoted by $x_{1}$ and $x_{2}$, are factorizable if and only if the corresponding state is separable. It indicates that the non-factorizability of the considered Gaussian wave functions expressed in these coordinates is equivalent to the entanglement of the corresponding state. Let us also note that, admittedly, a point transformation of the original position coordinates $(x_{1},x_{2})$, such as a centre of mass and relative coordinates, allows the examined Gaussian wave function to be factored. However, this does not imply the separability of the corresponding state. In order to demonstrate the findings of our study, we have proposed and investigated a theoretical model of a two-particle mechanical system that allows expounding physical interpretation of the presented result. The Hamiltonian we have derived describe a system consisting of two coupled anisotropic harmonic oscillators depending on the parameter $\alpha$. We have demonstrated that the limit form that corresponds to $\alpha\rightarrow 1_{-}$ or alternatively $\xi\to 0_{+}$ of the aforementioned Hamiltonian reduces to the Hamiltonian of the two independent anisotropic harmonic oscillators shifted in positions and momenta. Furthermore, we also have shown that assuming no shifts and assuming identical oscillation frequencies, the Hamiltonian we have obtained is consistent with the simplified model discussed in the previous work. Based on the results discussed in this paper, it can also be concluded that performing a unitary transformation corresponding to a canonical transformation on phase-space variables can lead to the decoupling of a system of interacting isotropic harmonic oscillators. However, the result obtained in this way leads to the loss of the individual characteristics of the input subsystems. Furthermore, each of the subsystems obtained in this way is based on new phase-space variables that mix the original variables used to describe the input systems.
@@ Appendix A Appendix: Mehler formulae
$\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}\frac{s^{m}t^{n}}{m!n!}H_{m,n}(z_{1},_{2})=\exp(sz_{1}+tz_{2}-st),$ (88)
@@
$=$ $e^{-\frac{1}{4}(\partial_{u}^{2}+\partial_{v}^{2})}\exp{\left[\sqrt{2}usz_{1}+\sqrt{2}vtz_{2}-2usvt\right]}.$
@@
Acknowledgements
The research project is partially supported by a subsidy from the Polish Ministry of Science and Higher Education and the program ’Initiative for Excellence – Research University’ for the AGH University of Krakow.